. One such theory hypothesizes that people should spontaneously follow a

24-hour cycle of sleeping and waking-even if they are not exposed to the usual pattern of

sunlight. To test this notion, 8 paid volunteers were placed in a room in which there was no light from the outside and no clocks or other indicators of time. After a month in the room, each individual developed a cycle. Their cycles at the end of the study were as follows: 25, 27, 25, 23, 24, 25, 26, 25. Using the .05 significance level, determine if these participants’ cycles are different from a 24-hour cycle.

SS = 10,

Sample mean = 25

1.what is the null hypothesis?

2.what is the research hypothesis?

3.What is the comparison mean?

4.what is the comparison distribution?

5.What is/are the cutoff(s)?

6.what is the t score for the sample score?

7.How is the denominator of a t-score different from the denominator of a z-score?

8.what is your conclusion for this hypothesis test?









1
Expert's answer
2021-11-01T20:00:52-0400

1.

The null hypothesis to be tested in this question is,

"H_0:\mu=24"

2.

The research hypothesis is the alternative hypothesis and it is given as,

"H_1:\mu\not=24"

3.

The comparison mean is the sample mean "\bar{x}" and its value is given as,

"\bar{x}=\sum(x)/n", where "n=8" and "\sum(x)=25+27+25+23+24+25+26+25=200"

Therefore,

"\bar{x}=200/8=25"

4.

The comparison distribution used is the student's t distribution. It is used because the population variance"(\mu)" is unknown and the sample size "(n=8)" is considerably small.

5.

The cutoffs are the t-table values that leaves an area of "\alpha/2=0.05/2=0.025" each to the right and to the left with "n-1=8-1=7" degrees of freedom. The values are given as,

"t_{0.025,7}=2.365" and "-t_{0.025,7}=-2.365" since the t distribution is symmetric about the mean.

Therefore, the cut off points are, "(-2.365,\space 2.365)"

6.

The t-score for the sample is given as,

"t^*=(\bar{x}-\mu)/(s/\sqrt{n})"

In order to determine the sample standard deviation "(s)" , we first evaluate for the variance given as,

"V(x)=\sum(x-\bar{x})^2/(n-1)=10/7=1.428571429"

Now,

"s=\sqrt{V(x)}=\sqrt{1.428571429}=1.1952(4\space dp)"

Thus "t^*=(25-24)/(1.1952/\sqrt{8})=1/0.422577127=2.3664(4\space dp)"

The t score for the sample score is "t^*=2.3664"

7.

The denominator of both scores is the standard error and it is given as,

"\sigma/\sqrt{n}" for a Z-score and "s/\sqrt{n}" for a t-score. The main difference is that the standard error "(s/\sqrt{n})" for a t-score is smaller than the standard error"(\sigma/\sqrt{n})" for a Z-score.

8.

For this hypothesis test, the null hypothesis is rejected if "|t^*|\gt t_{0.025,7}"

Since "|t^*|=|2.3664|=2.3664\gt t_{0.025,7}=2.365", we reject the null hypothesis and conclude that sufficient evidence exist to show that  participants’ cycles are different from a 24-hour cycle at 5% level of significance.